how can you figure out two symmetrical points from these two equations. y=5(x+6)^2 and y= -3/5(x-3)^2 I need to have the work done so i can learn from this and figure out other equations. I am not understanding. thank you.
Found 2 solutions by stanbon, dabanfield:
Answer by stanbon(75887)

how can you figure out two symmetrical points from these two equations. y=5(x+6)^2 and y= -3/5(x-3)^2 I need to have the work done so i can learn from this and figure out other equations. I am not understanding. thank you.
---
1st Equation:
y = 4(x+6)^2
Axis of symmetry is x = -6
--
Two points on the parabola that are symmetric to x= -6
would be:
Consider points 2 to the left of x=-6 and 2 to the right of -6.
Let x = -8, then y = 4(-8+6)^2 = 4*4 = 16, giving you point (-8,16)
Let x = -4, then y = 4(-4+6)^2 = 4*4 = 16, giving you point (-4,16)
======================================
2nd Equation:
y = (-3/5)(x-3)^2
Axia of symmetry is x = 3
Consides point 3 to the left and 3 to the right of x = 3
Let x = 0, then y = (-3/5)(-3)^2 = (-3/5)*9 = -27/5 giving you point(0,-27/5)
Let x = 6, then y = (-3/5)(6-3)^2 = (-3/5)9 = -27/5 giving you point(6,-27/5)
=============================================================================
Cheers,
Stan H.






Answer by dabanfield(803)

how can you figure out two symmetrical points from these two equations. y=5(x+6)^2 and y= -3/5(x-3)^2 I need to have the work done so i can learn from this and figure out other equations. I am not understanding. thank you.
Is the second equation y = -3/[(5*(x-3)^2] or y = (-3/5)*(x-3)^2?
Please clarify and resubmit. Also, by symmetrical points do you mean common simultaneous solutions to both equations?
Thanks.